OpenStudy (anonymous):

how do you do this? solve the equation for s v=s+12t over 9

OpenStudy (saifoo.khan):

\[v = \frac{s+12t}{9}\]This right>

OpenStudy (saifoo.khan):

\[9v = 12t + s\] \[s = 9v - 12t\]

jimthompson5910 (jim_thompson5910):

v=(s+12t)/9 9v=s+12t 9v-12t=s s=9v-12t

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np